OpenStudy (waheguru):

y-2=0 find teh slope and y - intercept HURRY HELP MEMMMEEEEE

OpenStudy (earthcitizen):

c=2, m=0

OpenStudy (waheguru):

yea can u explain!

OpenStudy (earthcitizen):

y=mx+c

OpenStudy (waheguru):

why is teh slope 0

OpenStudy (anonymous):

y=mx+b

OpenStudy (earthcitizen):

\[y=mx+c\]

hero (hero):

y - 2 = 0 y = 2 y = 2 is a horizontal line with no slope

OpenStudy (waheguru):

oh yea cuz a zero is NOTHING THANK Y U

OpenStudy (earthcitizen):

\[y-2=0 =>y=2=> y=0x+2\]

OpenStudy (earthcitizen):

no p!